OpenStudy (anonymous):

What is the value of x? when one angle of a triangle is 39 and the other is 102

OpenStudy (faiqraees):

x +39 +102 = 180 (Sum of all angles in a triangle is 180)

OpenStudy (anonymous):

ok so first ill add 39 and 102 together?

OpenStudy (simplymarie_x):

Correct, you'll add those together.

OpenStudy (anonymous):

39 and 102 = 141?

OpenStudy (simplymarie_x):

Correct. Now whats 180 - 141.

OpenStudy (anonymous):

so 39

OpenStudy (simplymarie_x):

Or if it's easier x + 141 = 180. You have to get x by itself. And yes that's right ^~^